# LLM with Relation Classifier for Document-Level Relation Extraction

### v1

@@ 5.1 Experimental Settings
Configuration   In the RCP stage, RoBERTalarge Liu et al. (2019) is chosen as the foundational PLM encoder. We implement early stopping based on the $F_{1}$ score obtained from the development set. We adopt AdamW Loshchilov and Hutter (2019) as the optimizer and apply a linear warmup for the learning rate at the first 6% steps. Main parameters are provided in Table 5. We use development set to manually tune the optimal parameters for the RCP stage based on the F1 score. The value of parameters we finally adopted are in bold.
@@
In the RC stage, we fine-tune LLMs with the RC-specific prompt, which is detailed in 4.2, using LoRA. Details regarding parameters are provided in Table 6. We train and evaluate the pre-classifier on a single NVIDIA A100 GPU during the RCP stage, and fine-tune LLMs on 4 NVIDIA A100 GPUs during the RC stage.
Evaluation   Consistent with other SOTA methods, we employ the standard evaluation metrics: $F_{1}$ and Ign $F_{1}$. Ign $F_{1}$ is calculated by excluding triplets that are already present in the training set from both the development and test sets.
5.2 Main Results
Method Dev Test
Ign $F_{1}$ $F_{1}$ Ign $F_{1}$ $F_{1}$
BERT-based
HIN-BERTbaseTang et al. (2020) 54.29 56.31 53.70 55.60
CorefBERTbaseYe et al. (2020) 55.32 57.51 54.54 56.96
CorefRoBERTalargeYe et al. (2020) 57.35 59.43 57.90 60.25
HeterGSAN+reconstructionXu et al. (2021b) 58.13 60.18 57.12 59.45
SSAN-RoBERTalargeXu et al. (2021a) 60.25 62.08 59.47 61.42
KD-RoBERTalargeTan et al. (2022a) 65.27 67.12 65.24 67.28
DREEAM-RoBERTalargeMa et al. (2023) 65.52 67.41 65.47 67.53
LLM-based
CoRSun et al. (2024) - 38.4 ± 10.6 - 38.5 ± 9.1
GenRDKSun et al. (2024) - 42.5 ± 10.6 - 41.5 ± 8.7
Our Methods
LoRA FT LLaMA2-7B-Chat 33.95 34.32 33.99 34.34
LoRA FT LLaMA2-13B-Chat 38.62 39.25 38.09 38.66
LMRC-LLaMA2-7B-Chat 52.40 54.10 52.81 54.73
LMRC-LLaMA2-13B-Chat 58.16 59.97 58.49 60.52
(a) Results on the development and test set of DocRED.
Method Ign $F_{1}$ $F_{1}$
BERT-based
KD-RoBERTalargeTan et al. (2022a) 77.60 78.28
DREEAMMa et al. (2023) 79.66 80.73
LLM-based
CoRSun et al. (2024) - 37.1 ± 9.2
GenRDKSun et al. (2024) - 41.3 ± 8.9
AutoRELilong et al. (2024) - 51.91
Our Methods
LoRA FT LLaMA2-7B-Chat 52.74 53.02
LoRA FT LLaMA2-13B-Chat 52.15 52.45
LMRC-LLaMA2-7B-Chat 72.33 72.92
LMRC-LLaMA2-13B-Chat 74.08 74.63
(b) Results on the test set of Re-DocRED
[…]
Overall Performance   As demonstrated in Table 7(a), the direct fine-tuning approach applied to LLaMA2 and other LLM-based methods expose suboptimal performance and inefficiencies in processing capabilities, underscoring the inherent challenges faced when leveraging LLMs for DocRE. Our results also corroborate the findings of Lilong et al. (2024). However, following the implementation of our proposed task division strategy, we observe a substantial improvement, marked by a significant boost in the $F_{1}$ score for LLaMA2 at both 7B and 13B scales. Table 7(b) presents a comparative analysis of LMRC’s performance against existing methods evaluated on the Re-DocRED test set. We observe that LMRC surpasses other LLM-based methods. Moreover, LMRC narrows the performance gap with the current SOTA method, DREEAM, positioning it as a promising paradigm for future DocRE.
Method Intra Inter
[…]
LoRA FT LLaMA2-13B-Chat 45.43 31.67
LMRC + LLaMA2-13B-Chat 65.88 52.66
[…]
Performance of Intra-/Inter-Extraction   We also report Intra-$F_{1}$/Inter-$F_{1}$, which respectively evaluate the model’s capability in capturing intra- and inter-sentence relations. LSR Nan et al. (2020) and GAIN Zeng et al. (2020), both graph-centric methodologies that leverages graph neural networks for inference, serve as valuable benchmarks. As illustrated in Table 8, LMRC not only outperforms the baseline achieved by directly fine-tuning LLaMA2 in both Intra- and Inter-$F_{1}$, but also maintains competitiveness with advanced graph-based models like GAIN-BERTbase. This demonstrates LMRC’s effectiveness, particularly in scenarios that require complex relational parsing from document-level context.
5.3 Fine-grained Analysis for Each Relation
To comprehensively assess the model’s sensitivity to different categories within the dataset, we undertake a fine-grained statistical analysis of the $F_{1}$ score for each relation type in the DocRED development set, comparing both the baseline and LMRC. Figure 3 visually showcases the enhancement effects in a more intuitive manner, highlighting the gains achieved by LMRC.
Our meticulous data analysis leads to the following conclusions: (1) LMRC enhanced the extraction performance for the majority of relations at both 7B and 13B scales. (2) A discernible positive relationship emerges between the frequency of occurrence of relations in the training set and their ultimate extraction performance. High-frequency relations, such as P569:"date of birth", achieved a remarkable $F_{1}$ score exceeding 90, showcasing the model’s proficiency in handling prevalent patterns.
However, it is noteworthy that despite these success, LMRC encounters challenges in extracting instances pertained to certain relations (e.g., P1336:"territory claimed by", P39:"position held"). We hypothesize that this limitation stems primarily from the scarcity of these relations in the training data, which restricts the model’s capacity to sufficiently learn the relevant features and nuances. Consequently, the model struggles to generalize effectively for these less frequently encountered relation types.
5.4 Ablation Studies
After showing the effectiveness of our proposed LMRC, we further explore the roles of RCP and RC stages on the DocRED development set. To validate the RCP stage, we fine-tune LLaMA2-13B-Chat to replace the original pre-classifier for binary classification. Moving on to the RC stage, we introduce a more simplified setup where we feed the model with entity pairs annotated in the ground truth. By masking relation tags, we challenge the model to classify these pairs into the predefined relation set. This process involves task-specific fine-tuning of LLaMA2-13B-Chat, specifically tailored for relation classification.
As shown in Table 9, substituting our pre-classifier with the fine-tuned LLaMA2-13B-Chat leads to a significant decrease in the $F_{1}$ score. We contend that the intricate multi-relation nature as well as the discrete distribution of relations in the DocRE task pose distinct challenges for LLMs, highlighting the crucial role of the pre-classifier in ensuring robust performance. Furthermore, the ablation study conducted on the RC stage underscores the proficiency of the fine-tuned LLaMA2-13B-Chat in performing relation classification. This achievement provides a solid foundation for future advancements, demonstrating the potential of LLMs excelling in complex relation extraction tasks.
Settings $F_{1}$ of RCP Ign $F_{1}$ $F_{1}$
RCP stage
LMRC 64.64 58.16 59.97
w/o pre-classifier w LLM 31.30 23.22 24.59
RC stage
relation classification - 86.09 86.75
Table 9: Ablation studies evaluated on DocRED dev set.
5.5 Out-of-Domain Relation Studies
In the aforementioned evaluation, we conservatively categorize all relations generated by LLaMA2 that do not align with the predefined relation set as erroneous outcomes1. However, previous research Wadhwa et al. (2023) has indicated that evaluating LLM-based models should not entirely rely on exact matches to targets. This is because LLMs can generate outputs that are semantically proximate to the target relations but may not satisfy the stringent exact-matching requirement. For instance, while the phrase "works at" in the model’s output resembles "work for" in the target, strict evaluation criteria would count it as a misclassification.
To gain a comprehensive understanding of this phenomenon, we delve deeper into the out-of-domain relations generated by LLaMA2-13B-Chat. We leverage SBERT [^2] https://www.sbert.net Reimers and Gurevych (2019), a well recognized approach for semantic similarity, to align out-of-domain relations to the predefined relation set $R$. This process first computes the cosine similarity between each out-of-domain relation and the relations within $R$. For each $r_{i}$ in the out-of-domain set, we choose the relation in $R$ that exhibits the highest similarity score $s_{\text{max}}^{i}$ as the optimal match. Intuitively, some out-of-domain relations may not carry meaningful information due to the hallucination issue of LLMs, and blindly aligning them to $R$ is not appropriate. Therefore, we introduce a heuristic threshold $\theta$. Only when $s_{\text{max}}^{i}\geq\theta$, does the alignment processed; otherwise, the triplets containing $r_{i}$ are discarded. After alignment, we recalculate the $F_{1}$ score of our methods on the DocRED dev set.
(a) LoRA FT
(b) LMRC
Figure 4: The impact of threshold $\theta$ for cosine similarity on the F1 score. Both methods are conducted on the DocRED dev set.
Figure 4 illustrates that LLaMA2 directly fine-tuned on DocRED attains its peak $F_{1}$ score when threshold $\theta$ is set approximately at 0.55. However, despite incorporating semantic alignment, the out-of-domain relations generated by LLaMA2 fine-tuned within the LMRC framework remain incorrect. We analyze these relations in depth and conclude that they mainly suffer from the following two problems:
  • While some of the generated out-of-domain relations could ostensibly be mapped to similar relations within the set $R$, the classification result is incorrect.
Additionally, it is noteworthy that both methods generated a relatively low number of out-of-domain relations, 74 and 72, respectively. This limited occurrence helps explain why adjustments to the threshold $\theta$ has little impact on the overall $F_{1}$ score.
5.6 Performance of GPT-4-turbo
To ensure a thorough assessment, we expand our investigation to encompass a broader array of models, notably incorporating GPT-3-turbo-16k and GPT-4-turbo into our study. This inclusion was made to validate the adaptability and robustness of our approach to DocRE across diverse model architectures.
We first leveraged a 3-shot learning format for GPT mdoels to directly tackle DocRE, utilizing examples meticulously curated from the human-annotated DocRED dataset. These examples were carefully chosen to include a diverse range of entity pairs, capturing both relation-expressing and non-relation entity pairs, thereby mirroring the complexity and variability inherent in real-world documents. Then, we adopted the LMRC paradigm. In this setting, GPT models took over the role of LLaMA2 in the RC stage, augmented by the addition of 3-shot learning format to enrich contextual understanding.
Methods Ign $F_{1}$ $F_{1}$ Intra $F_{1}$ Inter $F_{1}$
GPT-3.5-turbo-16k + 3-shot 6.57 6.97 11.78 4.24
GPT-4-turbo + 3-shot 9.85 10.71 14.52 7.43
LMRC + GPT-3.5-turbo-16k 27.27 28.39 38.52 17.19
LMRC + GPT-4-turbo 34.84 36.20 44.08 27.30
LMRC + LLaMA2-13B-Chat (FT) 58.16 59.97 65.88 52.66
Considering budget constraints, our experimental analysis was conducted on a randomly selected subset of 100 documents from the DocRED development set. The preliminary results are shown in Table 10. Before implementing LMRC, the extraction performances of GPT models were quite low, while their overall F1 scores were significantly improved after integrating LMRC strategy. This reflects the inherent challenges posed by DocRE and also demonstrates the effectiveness of our proposed method on various models. In addition, even GPT-4-turbo failed to surpass the fine-tuned LLaMA2-13B-Chat through 3-shot in-context learning, indicating the necessity of fine-tuning LLMs on DocRE tasks.
6 Conclusion

### v2

@@ 5.1 Experimental Settings
Configuration   In the RCP stage, RoBERTalarge Liu et al. (2019) is chosen as the foundational PLM encoder. We implement early stopping based on the $F_{1}$ score obtained from the development set. We adopt AdamW Loshchilov and Hutter (2019) as the optimizer and apply a linear warmup for the learning rate at the first 6% steps. Main parameters are provided in Table 7 in Appendix B. We use development set to manually tune the optimal parameters for the RCP stage based on the F1 score. The value of parameters we finally adopted are in bold.
In the RC stage, we fine-tune LLMs with the RC-specific prompt, which is detailed in 4.2, using LoRA. Details regarding parameters are provided in Table 8 in Appendix B. We train and evaluate the pre-classifier on a single NVIDIA A100 GPU during the RCP stage, and fine-tune LLMs on 4 NVIDIA A100 GPUs during the RC stage.
Evaluation   Consistent with other SOTA methods, we apply the standard evaluation metrics: $F_{1}$ and Ign $F_{1}$. Ign $F_{1}$ is calculated by excluding triplets that are already present in the training set from both the development and test sets.
5.2 Main Results
Method Dev Test
Ign $F_{1}$ $F_{1}$ Ign $F_{1}$ $F_{1}$
BERT-based
HIN-BERTbaseTang et al. (2020) 54.29 56.31 53.70 55.60
CorefBERTbaseYe et al. (2020) 55.32 57.51 54.54 56.96
CorefRoBERTalargeYe et al. (2020) 57.35 59.43 57.90 60.25
HeterGSAN+reconstructionXu et al. (2021b) 58.13 60.18 57.12 59.45
SSAN-RoBERTalargeXu et al. (2021a) 60.25 62.08 59.47 61.42
KD-RoBERTalargeTan et al. (2022a) 65.27 67.12 65.24 67.28
DREEAM-RoBERTalargeMa et al. (2023) 65.52 67.41 65.47 67.53
LLM-based
0-shot GPT-4Zhu et al. (2024) - - - 15.58
MDP GPT-4Zhu et al. (2024) - - - 21.33
FT Flan-UL2Peng et al. (2023) - - - 54.50
REPLM GPT3.5Ozyurt et al. (2023) - 59.66 - -
Our Methods
LoRA FT LLaMA2-7B-Chat 33.95 34.32 33.99 34.34
LoRA FT LLaMA2-13B-Chat 38.62 39.25 38.09 38.66
LMRC LLaMA2-7B-Chat 52.40 54.10 52.81 54.73
LMRC LLaMA2-13B-Chat 58.16 59.97 58.49 60.52
LMRC LLaMA3-8B-Instruct 59.13 61.01 58.77 60.84
LMRC LLaMA3.1-8B-Instruct 59.22 61.08 58.62 60.69
(a) Results on the development and test set of DocRED.
Method Ign $F_{1}$ $F_{1}$
BERT-based
KD-RoBERTalargeTan et al. (2022a) 77.60 78.28
DREEAMMa et al. (2023) 79.66 80.73
LLM-based
0-shot GPT-4Zhu et al. (2024) - 14.21
MDP GPT-4Zhu et al. (2024) - 20.16
AutoRE-Vicuna-7BLilong et al. (2024) - 53.84
Our Methods
LoRA FT LLaMA2-7B-Chat 52.74 53.02
LoRA FT LLaMA2-13B-Chat 52.15 52.45
LMRC LLaMA2-7B-Chat 72.33 72.92
LMRC LLaMA2-13B-Chat 74.08 74.63
LMRC LLaMA3-8B-Instruct 72.21 72.80
LMRC LLaMA3.1-8B-Instruct 72.35 72.93
(b) Results on the test set of Re-DocRED
[…]
Overall Performance   As demonstrated in Table 3(a), the direct fine-tuning approach applied to LLaMA models and other LLM-based methods expose suboptimal performance and inefficiencies in processing capabilities, underscoring the inherent challenges faced when leveraging LLMs for DocRE. Our results also corroborate the findings of Lilong et al. (2024). However, following the implementation of our proposed task division strategy, we observe a substantial improvement, marked by a significant boost in the $F_{1}$ score. Table 3(b) presents a comparative analysis of LMRC’s performance against existing methods evaluated on the Re-DocRED test set. We observe that LMRC surpasses other LLM-based methods. Moreover, compared to the poor performance of recent LLM-based methods, our LMRC framework achieves higher scores across a broader range of evaluations, narrowing the performance gap with the current SOTA method, DREEAM, and positioning it as a promising paradigm for future DocRE.
Method Intra Inter
[…]
LoRA FT LLaMA2-7B-Chat 52.74 53.02
LoRA FT LLaMA2-13B-Chat 52.15 52.45
LMRC LLaMA2-7B-Chat 72.33 72.92
LMRC LLaMA2-13B-Chat 74.08 74.63
LMRC LLaMA3-8B-Instruct 72.21 72.80
LMRC LLaMA3.1-8B-Instruct 72.35 72.93
[…]
Performance of Intra-/Inter-Extraction   We also report Intra-$F_{1}$/Inter-$F_{1}$, which respectively evaluate the model’s capability in capturing intra- and inter-sentence relations. LSR Nan et al. (2020) and GAIN Zeng et al. (2020), both graph-centric methodologies that leverages graph neural networks for inference, serve as valuable benchmarks. As illustrated in Table 4, LMRC not only outperforms our baseline obtained by directly fine-tuning LLaMA models in both Intra- and Inter-$F_{1}$, but also remains competitive with advanced graph-based models like GAIN-BERTbase. This demonstrates LMRC’s effectiveness, particularly in scenarios that require complex relational parsing within document-level context.
5.3 Ablation Studies
After demonstrating the effectiveness of our proposed LMRC, we further explore the roles of RCP and RC stages on the DocRED development set. To validate the RCP stage, we fine-tune LLaMA2-13B-Chat to replace the original pre-classifier for binary classification. For the RC stage, we introduce a more simplified setup where we feed the model with entity pairs annotated in the ground truth. By masking relation tags, we challenge the model to classify these pairs into the predefined relation set, fine-tuning LLaMA2-13B-Chat specifically for relation classification.
As shown in Table 5, substituting our pre-classifier with the fine-tuned LLaMA2-13B-Chat significantly reduced the $F_{1}$ score of the RCP stage as well as the overall extraction performance. We argue that the intricate multi-relation nature and the discrete distribution of relations in DocRE tasks present unique challenges for LLMs, highlighting the crucial role of our pre-classifier in ensuring robust performance. Furthermore, the ablation study conducted on the RC stage highlights the effectiveness of fine-tuned LLaMA2-13B-Chat in performing relation classification. This phenomenon demonstrates the potential of LLMs to excel in complex relation extraction tasks.
Settings $F_{1}$ of RCP Ign $F_{1}$ $F_{1}$
RCP stage
LMRC 64.64 58.16 59.97
w/o pre-classifier w LLM 31.30 23.22 24.59
RC stage
relation classification - 86.09 86.75
Table 5: Ablation studies evaluated on DocRED dev set.
5.4 Fine-grained Analysis for Each Relation
To further assess the model’s sensitivity to different categories within the dataset, we undertakes a fine-grained statistical analysis of the $F_{1}$ score for each relation type in the DocRED development set, comparing both baseline and LMRC. Figure 3 visually showcases the enhancement effects in a more intuitive manner.
Our meticulous data analysis leads to the following conclusions: (1) LMRC improves the extraction performance for the majority of relations. (2) A discernible positive relationship emerges between the frequency of relation occurrence in the training set and their corresponding extraction performance. High-frequency relations, such as P569: date of birth, achieved an improved $F_{1}$ score exceeding 90, showcasing the model’s proficiency in handling prevalent patterns.
However, it is noteworthy that despite these positive results, LMRC encounters challenges in extracting instances pertained to certain relations (e.g., P39: position held, P737: influenced by). We consider that this limitation stems primarily from the scarcity of these relations in the training data, which restricts the model’s capacity to sufficiently learn relevant features and nuances. Consequently, the model struggles to generalize effectively for these less frequently encountered relation types.
5.5 Out-of-Domain Relation Studies
In the aforementioned evaluation, we conservatively categorize all relations generated by LLMs that do not align with the predefined relation set as erroneous outcomes. However, previous research Wadhwa et al. (2023) has indicated that evaluating LLM-based models should not entirely rely on exact matches to targets. This is because LLMs can generate outputs that are semantically proximate to the target relations but may not satisfy the stringent exact-matching requirement. For instance, while the phrase "works at" in the model’s output resembles "work for" in the target, strict evaluation criteria would count it as a misclassification.
To gain a comprehensive understanding of this phenomenon, we delve deeper into the out-of-domain relations generated by LLaMA2-13B-Chat. We leverage SBERT Reimers and Gurevych (2019), a well recognized approach for semantic similarity, to align out-of-domain relations to the predefined relation set $R$. This process first computes the cosine similarity between each out-of-domain relation and the relations within $R$. For each $r_{i}$ in the out-of-domain set, we choose the relation in $R$ that exhibits the highest similarity score $s_{\text{max}}^{i}$ as the optimal match. Intuitively, some out-of-domain relations may not carry meaningful information due to the hallucination issue of LLMs, and blindly aligning them to $R$ is not appropriate. Therefore, we introduce a heuristic threshold $\theta$. Only when $s_{\text{max}}^{i}\geq\theta$, does the alignment processed; otherwise, the triplets containing $r_{i}$ are discarded. After alignment, we recalculate the $F_{1}$ score of our methods on the DocRED development set. Results are provided in Figure 4.
Figure 4: The impact of threshold $\theta$ for cosine similarity on the $F_{1}$ score. Left: LoRA fine-tune LLaMA2-13B-Chat. Right: LMRC with LLaMA2-13B-Chat. Both methods are conducted on the DocRED dev set.
As is shown, LLaMA2-13B-Chat directly fine-tuned on DocRED attains its peak $F_{1}$ score when threshold $\theta$ is set approximately at 0.55. However, despite incorporating semantic alignment, the out-of-domain relations generated by LLaMA2-13B-Chat fine-tuned within the LMRC framework remain incorrect. We analyze these relations in depth and conclude that they mainly suffer from the following problems:
  • While some of the generated out-of-domain relations could ostensibly be mapped to similar relations within set $R$, the classification result is incorrect.
Additionally, it is noteworthy that both methods generated a relatively low number of out-of-domain relations, 74 and 72, respectively. This limited occurrence helps explain why adjustments to the threshold $\theta$ has little impact on the $F_{1}$ score.
5.6 Performance of GPT models
To ensure a thorough assessment, we expand our investigation to encompass a broader array of models, notably incorporating GPT-3-turbo-16k and GPT-4-turbo into our study. This inclusion is made to validate the adaptability and robustness of our approach to DocRE across diverse model architectures.
We first leverage a 3-shot learning format for GPT mdoels to directly tackle DocRE, utilizing examples meticulously curated from the human-annotated DocRED dataset. These examples are carefully chosen to include a diverse range of entity pairs, capturing both relation-expressing and non-relation entity pairs, thereby mirroring the complexity and variability inherent in real-world documents. Then, we adopt the LMRC paradigm. In this setting, GPT models take over the role of LLaMA models in the RC stage, augmented by the addition of 3-shot learning format to provide essential examples and enrich contextual understanding.
Methods Ign $F_{1}$ $F_{1}$ Intra $F_{1}$ Inter $F_{1}$
GPT-3.5-turbo-16k (3-shot) 6.57 6.97 11.78 4.24
GPT-4-turbo (3-shot) 9.85 10.71 14.52 7.43
LMRC + GPT-3.5-turbo-16k (3-shot) 27.27 28.39 38.52 17.19
LMRC + GPT-4-turbo (3-shot) 34.84 36.20 44.08 27.30
LMRC + LLaMA3.1-8B-Instruct (FT) 59.22 61.08 66.70 54.17
Considering budget constraints, our experimental analysis is conducted on a randomly selected subset of 100 documents from the DocRED development set. The preliminary results are shown in Table 6. Before implementing LMRC, the extraction performances of GPT models are quite low, while their overall F1 scores are significantly improved after integrating LMRC strategy. This reflects the inherent challenges posed by DocRE and also demonstrates the effectiveness of our proposed method on various models. In addition, even GPT-4-turbo failed to surpass the fine-tuned LLaMA3.1-8B-Instruct through 3-shot in-context learning, indicating the necessity of fine-tuning LLMs on DocRE tasks.
6 Conclusion
Appendix A Prompts
Prompt for Document-level Relation Extraction Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exists relations, select relations for the entity pairs from the relation set; if there is no relation, return None. The format of the input entity pair is ‘(head entity| -| tail entity)’. Your output format is ‘(head entity| relation/None| tail entity)’. ### Relation set: {predefined relation set} ### Text: {text} ### {number of entity pairs} Entity pairs: {entity pairs} ### Response:
Prompt for Relation Classification Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: This is a relation classification task. we will provide entity pairs that require relation classification. Your task is to select relations for each entity pair from the given relation set based on the information in the text. There may be multiple relations between an entity pair. The format of the input entity pair is ‘(head entity| -| tail entity)’. Your output format is ‘(head entity| relation| tail entity)’. ### Relation set: {predefined relation set} ### Text: {text} ### {number of entity pairs} Entity pairs: {entity pairs} ### Response:
Appendix B Parameter settings